OpenStudy (ineedhelp10):

it takes 1.25 seconds for light to travel from the moon to the earth. how many miles away is the moon?

OpenStudy (ineedhelp10):

@Directrix help

OpenStudy (anonymous):

Look it up on Google!

OpenStudy (anonymous):

jk

OpenStudy (ineedhelp10):

i tried that but it didnt make sense

OpenStudy (unklerhaukus):

Use the speed equals distance over time \[\text{speed} =\frac{\text{distance}}{\text{time}}\] And the speed of light in miles per second is \[c=1.863\times10^{5} [\text{mi}/\text{s}]\]

OpenStudy (ineedhelp10):

how did you get 1.863?

OpenStudy (unklerhaukus):

well, in SI \[c = 2.998\times10^8 [\text{m}/\text{s}]\] and a mile is \[[\text{mi}]=1.609\times10^3[\text {m}]\]

OpenStudy (unklerhaukus):

So \[c = 2.998\times10^8 [\text{m}/\text{s}]\times\frac{\text{mi}}{1.609\times10^3[\text m]}\\\quad=1.863\times10^5[\text{mi}/\text s]\]

OpenStudy (ineedhelp10):

okay so the answer would be 1.863 x 10 to the 5th power right

OpenStudy (unklerhaukus):

You need you calculate the distance to the moon (in miles), not the speed of light

OpenStudy (unklerhaukus):

\[c=d/t\] \[d = ct\]

OpenStudy (unklerhaukus):

So multiply the speed of light in [mi/s], by the time in [s]

OpenStudy (ineedhelp10):

so i would have to multiply 2.998 x 10 (eighth power) x the speed of light or multiply the speed of light by something else?

OpenStudy (unklerhaukus):

No, c = 2.998 x 10^8 [m/s] is the speed of light in SI we converted this to miles/second c = 1.863 x10^5 [mi/s] Now we have c in [mi/s] used the formula for speed to find the distance,

OpenStudy (ineedhelp10):

ooooh so i would multiply 1.863 x 10^5

OpenStudy (ineedhelp10):

another step i meant

OpenStudy (unklerhaukus):

\[d = ct\\ \quad=1.863\times10^5\,[\text{mi}/\text s]\times1.25\,[\text s]\]

OpenStudy (ineedhelp10):

okay thank you so much

OpenStudy (unklerhaukus):

So what is the distance to the moon?

OpenStudy (ineedhelp10):

when i did the equation i got 232,875 miles

OpenStudy (unklerhaukus):

good work! \(\checkmark \)

OpenStudy (ineedhelp10):

=)